These were accepted as a reasonable account of the last meeting.
It was agreed that the review of pages with a review date prior to mid-2014 should be completed by the beginning of April. (The first Ops meeting in April is Wed 13th.)
All kit to be kept must be labelled by 1st March.
If we have any volunteers, we should have a couple of server room tours. If you would like to volunteer for the LCFG workshop in FH on Wednesday 17th, contact Alastair.
First stage (inventory and cull) in progress. Please reply to Carol's email as soon as possible.
MPU, RAT (almost definitely) and Inf done.
Tracked at CEG - using wire 'O' for now.
User documentation complete but documentation for C(S)Os still required
Key points identified. To be fleshed out. The Computer Security course has some useful pointers.
Actions added or revived:
Infrastructure to look into this.
Comments/changes to Craig/Chris's notes (which can be found in COs AFS area) to be sent to COs by next Ops meeting. (Minor changes to go straight to Craig/Chris.) Thoughts on where to go from here to be sent to COs.
Create an 'Emergency' page on computing.help with details of what will happen if an emergency occurs.
We should contact Angus for advice.
Transferred to RAT to-do list so removed from Ops list.
Yes. Workaround in progress. On Services to-do list.
Tim/Alison discussed. Are confident that we have enough people with sufficient knowledge even if not physically present.
Done. User documentation
Report from Computing Executive Group
Reply to COs with comments/changes to notes produced by Craig/Chris along with suggestions about how to progress by next Operational Meeting. (If minor corrections, just contact Craig/Chris directly).
Just in case you weren't aware, kit in B.03 that you want to be kept must be labelled by end of Feb otherwise it will be recycled.
Reports from Units
Please give this a go and report any issues to the Inf Unit.
Please do not filter out/ignore important informational emails (e.g. DHCP errors). They are sent so that appropriate action can be taken.
Please note the list of server models whose headers have been fixed for the more modern naming convention. You can use LCFG_NETWORK_LEGACY_NAME to use the legacy-style naming convention.
Note that the default on develop will soon be SL7.2. Once ready, it will become the default for all desktops with the exception of lab machines (which will remain on 7.1 until the end of teaching. The student.login and student.compute will also be kept on 7.1 to match the labs. (They have still to be upgraded from SL6). We need a volunteer to check kde on SL7.2.
Please reply to Carol's email about KVM usage before next Operational Meeting.
While we work on getting DICE running on the G2s, we need to keep all remaining G1s for installing as DICE machine i.e. not MDP or self-managed.
A further 2TB of general AFS space has been created.
We are not making as much progress as we would like. This will be made a priority task with a target of beginning of May as the target date.
Topics for discussion
Here's a copy of the notes from the wiki page:
It was agreed that remote ssh access for COs is not the way to go.
Inf will look into running a nut service for SMSR.
Will put sign up on door with our shutdown procedure.
Will ensure procedure is documented on computing.help - an 'Emergency' page.
There is uncertainty over the legality of our installing an exFAT FUSE module on DICE. We should contact Angus for advice on licence issues.
The next meeting will be on 24th February 2016 at 10:00 in IF-4.31.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh